# Minimum Volume Conformal Sets for Multivariate Regression

### v1

@@ 1 Introduction
In predictive modeling, quantifying uncertainty is often as crucial as making accurate predictions. Traditional point estimates provide limited insight into predictive accuracy, whereas prediction sets offer a more robust alternative by identifying regions that contain the true outcome with high probability. Conformal prediction [59, 49, 2] provides a model-agnostic framework for constructing such sets with finite-sample validity, ensuring that the true response is captured at least $1-\alpha$ fraction of the time without requiring strong distributional assumptions.
In the setting of univariate regression, conformal prediction can produce prediction intervals that adapt to heteroskedasticity in the data. Quantile regression, optimized using the pinball loss, is a common approach for learning such intervals [45]. Extending these ideas to multivariate regression, however, where the response is vector-valued, introduces significant challenges. A straightforward extension—constructing Cartesian products of marginal intervals [37]—fails to account for dependencies across dimensions, resulting in overly conservative and inefficient prediction sets. Instead, prediction sets must be structured to adapt to the joint distribution of the residuals, balancing validity, efficiency, and flexibility.
More structured alternatives, such as ellipsoidal prediction sets shaped by empirical residual dependencies [23, 35], incorporate covariance information but assume elliptical symmetry, limiting their adaptability to more complex distributions. More flexible approaches attempt to model joint dependencies explicitly. Copula-based methods [34, 52] and density estimators [22, 60] relax geometric constraints but often require accurate joint distribution estimation, which is computationally expensive in high dimensions, and typically leads to high estimation variance. Other methods, including optimal transport [25, 55] and quantile region estimation [14], allow for nonconvex and multimodal structures, improving adaptability. However, these approaches typically lack an explicit volume minimization criterion, leading to unnecessarily large sets, and often involve computationally expensive procedures that scale poorly in high dimensions. These methods, along with other approaches, will be reviewed in more detail in Section 1.1 on related work.
@@
Figure 1: MVCSs obtained using the first-order optimization strategy described in Section 2.2 with coverage level $0.95$. The left figure is a single-norm MVCS fit with $p=1.98$ for Gaussian data, the middle one is a single-norm MVCS fit with $p=0.94$ for exponential data, while the right one is a multi-norm MVCS optimized over four different norms, yielding region-specific values of $p=0.50,0.49,2.52,1.11$, for a combination of a uniform distribution over three unit balls with norms $0.5$, $3$ and $1$.
@@ 1.1 Related work
Several approaches have been proposed to construct valid prediction sets while balancing flexibility and efficiency [10]. We review key contributions in this area, highlighting their strengths and limitations.
@@
Hyperrectangles. One of the earliest approaches constructs prediction sets as Cartesian products of marginal intervals, resulting in hyperrectangles [37]. While straightforward and computationally efficient, this method fails to capture dependencies among response variables, often leading to overly conservative prediction regions.
@@
Ellipsoidal prediction sets. Ellipsoidal prediction sets provide a more structured alternative, using empirical covariance matrices to shape prediction regions. The work [23] introduced a method that constructs a single, global ellipsoidal region using the sample covariance matrix. Building on this, [35] incorporated k-nearest neighbors to estimate local covariance, blending it with the global covariance matrix to improve adaptability to local structure. Finally, [19] further extended these methods by developing a rigorous mathematical framework, establishing theoretical guarantees, and leveraging empirical covariance matrices for calibration and efficiency. However, while these approaches can yield tighter prediction sets in some cases, they remain restricted to convex shapes and assume an underlying elliptical structure, limiting their flexibility in capturing more complex distributions.
@@
Convex templates. To address multimodality, [58] introduced a clustering-based method that fits separate convex templates—such as convex hulls, hyperrectangles, and ellipsoids—to each cluster. This improves flexibility in capturing diverse residual distributions while still maintaining convexity. Additionally, although not directly focused on prediction sets, related work has studied minimum-volume enclosing polytopes [39, 48].
@@
Copula-based approaches. Rather than relying on fixed geometric assumptions, copula-based methods explicitly model dependencies in the response distribution. The works [34] and [52] leveraged copulas to construct (joint) prediction sets, offering a more data-driven approach that avoids strong parametric assumptions. These methods are inherently adaptive, as they adjust prediction regions based on the estimated dependence structure of the data. However, their effectiveness depends on accurate copula estimation, which can be challenging in high dimensions.
@@
Density-based and sampling-based approaches. Another approach to constructing prediction sets involves directly estimating the conditional distribution of the response. In this direction, [22] used conditional density estimation to derive predictive sets, leveraging density estimators to approximate the distribution of residuals. Alternatively, [60] employed generative models to sample from the predictive distribution, capturing complex dependencies that density estimation alone might not fully capture. Moreover, [42] introduced a hybrid framework that combines both density estimation and generative modeling, improving the flexibility of predictive set construction, particularly in cases with high heteroskedasticity. A related line of work considers density-based quantile methods, where quantile regions are estimated using conditional density level sets [6].
@@
Latent-space quantile methods. Nonconvex prediction sets can be constructed by mapping conditional distributions into a latent space where level sets remain convex, then transforming them back into the original space. This approach, proposed by [14], combines directional quantile regression with conditional variational autoencoders to effectively capture complex distributional features.
@@
Optimal transport-based methods. Optimal transport has been used to construct multivariate conformal prediction sets by defining meaningful rankings in multidimensional spaces. In this direction, [55] built upon the developments in [8, 17] to map multivariate residuals onto a uniform distribution using transport maps, enabling a structured approach to defining multivariate nonconformity scores. Concomitantly, [25] further improved computational efficiency by leveraging differentiable transport maps and entropic regularization techniques, facilitating scalable nonconformity score computations. While these methods provide a principled approach to multivariate prediction regions, they require solving transport optimization problems, which are computationally demanding in high dimensions. Finally, transport-based quantile methods, such as vector quantile regression [7] and center-outward quantiles [16], leverage optimal transport maps to define multivariate quantile regions. While these methods do not construct conformal prediction sets, they provide an alternative geometric approach to ordering multivariate data and characterizing uncertainty.
@@
Extensions to functional and distributional data. Conformal prediction has also been extended beyond finite-dimensional regression to handle more complex data structures. One major direction is functional data, where [31] introduced a conformal framework, later refined by [12], to improve adaptivity for structured functional domains. In the context of distributional regression, [9] developed conformal methods that provide coverage guarantees in probability spaces. More broadly, [27] formulated a general conformal prediction framework for arbitrary metric spaces, expanding its applicability to a wide range of structured data types.
@@
Volume-minimizing methods. Among prior works, [19] and [58] are the most closely related to our method, as both focus explicitly on minimizing the volume of the prediction set. However, [58] restricts prediction regions to convex shapes and employs heuristic clustering algorithms to adaptively partition the data, while [19] optimizes volume under the assumption of elliptical distributions, relying on empirical covariance estimation. Additionally, although not directly focused on prediction sets, a closely related line of research has investigated the problem of finding the minimum-volume enclosing ellipsoid for a given set of points [51, 56, 57, 53, 5, 46, 29, 18, 36]. Our method generalizes these approaches by optimizing prediction regions over arbitrary norms, removing restrictive convexity assumptions while maintaining exact finite-sample coverage.
@@
Local adaptivity. Beyond their structural differences, some of these methods are designed to be adaptive, adjusting prediction sets based on the underlying data distribution [55, 35, 12, 9, 28, 33, 34, 27, 22, 60, 43, 58, 14, 10], while others provide theoretical guarantees of asymptotic conditional coverage [55, 9, 22, 43]. Our method belongs to the class of adaptive approaches.
@@ 1.2 Preliminaries on conformal prediction
We focus on split conformal prediction, one of the most widely used methods for constructing valid prediction sets due to its simplicity and computational efficiency [40, 32]. Given a dataset $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{N}\sim\mathbb{P}$, where each $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ represents a covariate vector, with $d$ components that we refer to as features, and each $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}^{k}$ represents a multivariate response, the goal is to construct a prediction region $C(X)$ for a new test covariate $X$ such that it contains the true response $Y$ with probability at least $1-\alpha$.
@@
[…]
@@
Given a set $\{a_{i}\}_{i=1}^{n}\subset\mathbb{R}$, we denote by $\sigma_{r}\{a_{i}\}$ the $r$-th largest element of the set, and by $\overline{\sigma}_{r}\{z_{i}\}$ the average of the $r$ largest values of the set, i.e., $\overline{\sigma}_{r}\{z_{i}\}:=\frac{1}{r}\sum_{j=1}^{r}\sigma_{j}\{z_{i}\}$. We denote by $\mathds{1}\{x\in A\}$ the indicator function, equal to $1$ if $x\in A$ and to 00 if $x\notin A$, for some set $A$. We denote by $\mathrm{SO}(k)$ the group of rotation matrices in $\mathbb{R}^{k\times k}$, given by $\mathrm{SO}(k)\coloneqq\{R\in\mathbb{R}^{k\times k}\mid R^{\top}R=I,\;\det(R)=1\}$. Moreover, we denote by $\mathrm{Diag}(k)$ the space of diagonal matrices in $\mathbb{R}^{k\times k}$. We assume that if the objective function of a minimization (maximization) problem is expressed as the difference of two terms, both evaluating to $\infty$, then the objective function value is interpreted as $\infty$ ($\infty$- ∞). This aligns with the extended arithmetic rules in [44].
@@ 2 Minimum-Volume Covering Set
[…]
@@
[…]
@@ 2.1 Arbitrary norm-based sets
[…]
@@
Here, the constraint ensures that the set contains at least $n-r+1$ of the given points $\{y_{1},\dots,y_{n}\}$. The quantity $r$ is chosen such that this corresponds to a fraction $1-\alpha$ of the total points. The volume of a norm-based set is given by
@@
[…]
@@
Throughout this section, we refer to problem (5) as a minimum-volume covering set (MVCS) problem. While the objective function $-\log\det(M)$ is convex for $M\succcurlyeq 0$, the problem remains nonconvex due to the cardinality constraint, which introduces a combinatorial structure to the feasible set. This problem is known to be NP hard (see [1, Proposition 1]). The following proposition, based on a homogeneity argument, provides an exact reformulation that enables both convex relaxations and exact reformulations via difference-of-convex (DC) programming. In Section 2.2 we also formulate this loss in a way that allows the use of first-order optimization strategies.
@@
Now, due the cardinality constraint, for given $\Lambda$ and $\eta$, the optimal $\nu^{\star}(\Lambda,\eta)$ is equal to the $(n-r+1)$-th smallest value in the set $\left\{\|\Lambda y_{i}+\eta\|\right\}_{i=1}^{n}$, or, equivalently, the $r$-th largest value, i.e.,
@@
$\begin{split}\min\quad&-\log\det(\zeta\Lambda)+\sigma_{r}\left\{\|\zeta(\Lambda y_{i}+\eta)\|\right\}\\ \mathrm{s.t.}\quad&\Lambda\succcurlyeq 0,\;\eta\in\mathbb{R}^{k},\zeta\geq 0,\\ \end{split}$
whose objective function can be rewritten as $-k\log\zeta-\log\det(\Lambda)+\zeta\sigma_{r}\left\{\|\Lambda y_{i}+\eta\|\right\}$. Now, since the objective function is convex and differentiable in $\zeta$, we can minimize explicitly over $\zeta$ and obtain
@@
where $\overline{\sigma}_{r}$ and $\overline{\sigma}_{r-1}$ denote the averages of the $r$ and $r-1$ largest values in the set. A formal proof that $f$ and $g$ are convex functions is provided in Appendix A.
A key advantage of the DC reformulation in (9) is that it allows for efficient local optimization using the DC algorithm (DCA) [54, 20]. This method iteratively constructs convex approximations of the objective function, solving a convex subproblem at each step rather than optimizing the original nonconvex function directly. This approach ensures scalability while preserving the problem’s structure. We briefly outline the DCA approach.
@@
[…]
@@
In the special case of $p=2$, this reduces to $g_{i}={\Lambda_{t}y_{i}+\eta_{t}}/{\|\Lambda_{t}y_{i}+\eta_{t}\|_{2}}$. In practice, problem (10) can be solved efficiently using cvxpy [11] with the Mosek solver [3].
A key property of the DCA is that the sequence of objective values $f(\Lambda_{t},\eta_{t})-g(\Lambda_{t},\eta_{t})$ is nonincreasing and convergent (see, e.g., [38, Lemma 1]). However, since problem (6) is nonconvex, DCA does not guarantee global optimality. The algorithm may converge to a local minimum, meaning that the resulting prediction set may not have the smallest possible volume.
@@
Let $(\Lambda^{\star},\eta^{\star})$ denote the final iterate of the DCA method (10) or the global solution of the convex relaxation (11). The corresponding MVCS set, which contains $n-r+1$ points, is given by $\{y\in\mathbb{R}^{k}\mid\|\Lambda^{\star}y+\eta^{\star}\|\leq\sigma_{r}\left\{\|\Lambda^{\star}y_{i}+\eta^{\star}\|\right\}\}$. We illustrate in Figure 2 the solutions obtained by both methods for different distributions. In the Gaussian case (left), the solutions are nearly identical, with both approaches aligning with the empirical covariance structure. However, for an asymmetric distribution (right), the convex relaxation method is more influenced by the tails, stretching the prediction set outward, whereas the DCA solution remains concentrated around the high-density region near the origin. These observations suggest that in elliptical distributions, the convex relaxation method performs well, but for non-elliptical distributions, the DCA method provides tighter, more representative sets.
@@
A key insight emerges when taking the derivative of the objective function in (11) with respect to $\nu$: at optimality, $\nu^{\star}$ is the $(n-r+1)$-th order statistic of the nonconformity scores, enforcing the required coverage constraint. This mechanism closely parallels the structure of quantile regression [26], where the pinball loss implicitly selects a quantile level by balancing over- and under-estimation errors. A similar argument appears in [47, Lemma 2.2.1], which establishes that minimizing the pinball loss aligns with quantile estimation through a first-order optimality condition.

### v2

@@ 1 Introduction
In predictive modeling, quantifying uncertainty is often as crucial as making accurate predictions. Traditional point estimates provide limited insight into predictive accuracy, whereas prediction sets offer a more robust alternative by identifying regions that contain the true outcome with high probability. Conformal prediction Vovk et al. [2005]; Shafer and Vovk [2008]; Angelopoulos et al. [2024] provides a model-agnostic framework for constructing such sets with finite-sample validity, ensuring that the true response is captured at least $1-\alpha$ fraction of the time without requiring strong distributional assumptions.
In the setting of univariate regression, conformal prediction can produce prediction intervals that adapt to heteroskedasticity in the data. Quantile regression, optimized using the pinball loss, is a common approach for learning such intervals Romano et al. [2019]. Extending these ideas to multivariate regression, however, where the response is vector-valued, introduces significant challenges. A straightforward extension—constructing Cartesian products of marginal intervals Neeven and Smirnov [2018]—fails to account for dependencies across dimensions, resulting in overly conservative and inefficient prediction sets. Instead, prediction sets must be structured to adapt to the joint distribution of the residuals, balancing validity, efficiency, and flexibility.
More structured alternatives, such as ellipsoidal prediction sets shaped by empirical residual dependencies Johnstone and Cox [2021]; Messoudi et al. [2022], incorporate covariance information but assume elliptical symmetry, limiting their adaptability to more complex distributions. More flexible approaches attempt to model joint dependencies explicitly. Copula-based methods Messoudi et al. [2021]; Sun and Yu [2022] and density estimators Izbicki et al. [2022]; Wang et al. [2023] relax geometric constraints but often require accurate joint distribution estimation, which is computationally expensive in high dimensions, and typically leads to high estimation variance. Other methods, including optimal transport Klein et al. [2025]; Thurin et al. [2025] and quantile region estimation Feldman et al. [2023], allow for nonconvex and multimodal structures, improving adaptability. However, these approaches typically lack an explicit volume minimization criterion, leading to unnecessarily large sets, and often involve computationally expensive procedures that scale poorly in high dimensions. These methods, along with other approaches, will be reviewed in more detail in Section 1.1 on related work.
@@
Figure 1: MVCSs obtained using the first-order optimization strategy described in Section 2.2 with coverage level $0.95$. The left figure is a single-norm MVCS fit with $p=1.98$ for Gaussian data, the middle one is a single-norm MVCS fit with $p=0.94$ for exponential data, while the right one is a multi-norm MVCS optimized over four different norms, yielding region-specific values of $p=0.50,0.49,2.52,1.11$, for a combination of a uniform distribution over three unit balls with norms $0.5$, 33 and $1$.
@@ 1.1 Related work
Several approaches have been proposed to construct valid prediction sets while balancing flexibility and efficiency Dheur et al. [2025]. We review key contributions in this area, highlighting their strengths and limitations.
@@
Hyperrectangles. One of the earliest approaches constructs prediction sets as Cartesian products of marginal intervals, resulting in hyperrectangles Neeven and Smirnov [2018]. While straightforward and computationally efficient, this method fails to capture dependencies among response variables, often leading to overly conservative prediction regions.
@@
Ellipsoidal prediction sets. Ellipsoidal prediction sets provide a more structured alternative, using empirical covariance matrices to shape prediction regions. The work Johnstone and Cox [2021] introduced a method that constructs a single, global ellipsoidal region using the sample covariance matrix. Building on this, Messoudi et al. [2022] incorporated k-nearest neighbors to estimate local covariance, blending it with the global covariance matrix to improve adaptability to local structure. Finally, Henderson et al. [2024] further extended these methods by developing a rigorous mathematical framework, establishing theoretical guarantees, and leveraging empirical covariance matrices for calibration and efficiency. However, while these approaches can yield tighter prediction sets in some cases, they remain restricted to convex shapes and assume an underlying elliptical structure, limiting their flexibility in capturing more complex distributions.
@@
Convex templates. To address multimodality, Tumu et al. [2024] introduced a clustering-based method that fits separate convex templates—such as convex hulls, hyperrectangles, and ellipsoids—to each cluster. This improves flexibility in capturing diverse residual distributions while still maintaining convexity. Additionally, although not directly focused on prediction sets, related work has studied minimum-volume enclosing polytopes Panigrahy [2004]; Saha et al. [2011].
@@
Copula-based approaches. Rather than relying on fixed geometric assumptions, copula-based methods explicitly model dependencies in the response distribution. The works Messoudi et al. [2021] and Sun and Yu [2022] leveraged copulas to construct (joint) prediction sets, offering a more data-driven approach that avoids strong parametric assumptions. These methods are inherently adaptive, as they adjust prediction regions based on the estimated dependence structure of the data. However, their effectiveness depends on accurate copula estimation, which can be challenging in high dimensions.
@@
Density-based and sampling-based approaches. Another approach to constructing prediction sets involves directly estimating the conditional distribution of the response. In this direction, Izbicki et al. [2022] used conditional density estimation to derive predictive sets, leveraging density estimators to approximate the distribution of residuals. Alternatively, Wang et al. [2023] employed generative models to sample from the predictive distribution, capturing complex dependencies that density estimation alone might not fully capture. Moreover, Plassier et al. [2024a] introduced a hybrid framework that combines both density estimation and generative modeling, improving the flexibility of predictive set construction, particularly in cases with high heteroskedasticity. A related line of work considers density-based quantile methods, where quantile regions are estimated using conditional density level sets Camehl et al. [2024].
@@
Latent-space quantile methods. Nonconvex prediction sets can be constructed by mapping conditional distributions into a latent space where level sets remain convex, then transforming them back into the original space. This approach, proposed by Feldman et al. [2023], combines directional quantile regression with conditional variational autoencoders to effectively capture complex distributional features.
@@
Optimal transport-based methods. Optimal transport has been used to construct multivariate conformal prediction sets by defining meaningful rankings in multidimensional spaces. In this direction, Thurin et al. [2025] built upon the developments in Chernozhukov et al. [2017]; Hallin et al. [2021] to map multivariate residuals onto a uniform distribution using transport maps, enabling a structured approach to defining multivariate nonconformity scores. Concomitantly, Klein et al. [2025] further improved computational efficiency by leveraging differentiable transport maps and entropic regularization techniques, facilitating scalable nonconformity score computations. While these methods provide a principled approach to multivariate prediction regions, they require solving transport optimization problems, which are computationally demanding in high dimensions. Finally, transport-based quantile methods, such as vector quantile regression Carlier et al. [2016] and center-outward quantiles del Barrio et al. [2024], leverage optimal transport maps to define multivariate quantile regions. While these methods do not construct conformal prediction sets, they provide an alternative geometric approach to ordering multivariate data and characterizing uncertainty.
@@
Extensions to functional and distributional data. Conformal prediction has also been extended beyond finite-dimensional regression to handle more complex data structures. One major direction is functional data, where Lei et al. [2015] introduced a conformal framework, later refined by Diquigiovanni et al. [2021], to improve adaptivity for structured functional domains. In the context of distributional regression, Chernozhukov et al. [2021] developed conformal methods that provide coverage guarantees in probability spaces. More broadly, Kuchibhotla [2020] formulated a general conformal prediction framework for arbitrary metric spaces, expanding its applicability to a wide range of structured data types.
@@
Volume-minimizing methods. Among prior works, Henderson et al. [2024] and Tumu et al. [2024] are the most closely related to our method, as both focus explicitly on minimizing the volume of the prediction set. However, Tumu et al. [2024] restricts prediction regions to convex shapes and employs heuristic clustering algorithms to adaptively partition the data, while Henderson et al. [2024] optimizes volume under the assumption of elliptical distributions, relying on empirical covariance estimation. Additionally, although not directly focused on prediction sets, a closely related line of research has investigated the problem of finding the minimum-volume enclosing ellipsoid for a given set of points Sun and Freund [2004]; Todd [2016]; Todd and Yıldırım [2007]; Tao et al. [2017]; Bowman and Heath [2023]; Rosa and Harman [2022]; Kumar and Yildirim [2005]; Harris et al. [2024]; Mittal and Hanasusanto [2022]. Our method generalizes these approaches by optimizing prediction regions over arbitrary norms, removing restrictive convexity assumptions while maintaining exact finite-sample coverage.
@@
Local adaptivity. Beyond their structural differences, some of these methods are designed to be adaptive, adjusting prediction sets based on the underlying data distribution Thurin et al. [2025]; Messoudi et al. [2022]; Diquigiovanni et al. [2021]; Chernozhukov et al. [2021]; Kuleshov et al. [2018]; Messoudi et al. [2020, 2021]; Kuchibhotla [2020]; Izbicki et al. [2022]; Wang et al. [2023]; Plassier et al. [2024b]; Tumu et al. [2024]; Feldman et al. [2023]; Dheur et al. [2025], while others provide theoretical guarantees of asymptotic conditional coverage Thurin et al. [2025]; Chernozhukov et al. [2021]; Izbicki et al. [2022]; Plassier et al. [2024b]. Our method belongs to the class of adaptive approaches.
@@ 1.2 Preliminaries on conformal prediction
We focus on split conformal prediction, one of the most widely used methods for constructing valid prediction sets due to its simplicity and computational efficiency Papadopoulos et al. [2002]; Lei et al. [2018]. Given a dataset $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{N}\sim\mathbb{P}$, where each $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ represents a covariate vector, with $d$ components that we refer to as features, and each $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}^{k}$ represents a multivariate response, the goal is to construct a prediction region $C(X)$ for a new test covariate $X$ such that it contains the true response $Y$ with probability at least $1-\alpha$.
@@
A key component in achieving this guarantee is the choice of a nonconformity score function $s:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, which measures how atypical an observed response $Y$ is relative to the model’s prediction for a given covariate vector $X$. A standard choice in multivariate regression is the norm of the residual $s(X,Y)=\|f(X)-Y\|$, where $\|\cdot\|$ is typically the Euclidean norm, though other norms can be considered to better capture the structure of the residuals.
@@
Given a set $\{a_{i}\}_{i=1}^{n}\subset\mathbb{R}$, we denote by $\sigma_{r}\{a_{i}\}$ the $r$-th largest element of the set, and by $\overline{\sigma}_{r}\{z_{i}\}$ the average of the $r$ largest values of the set, i.e., $\overline{\sigma}_{r}\{z_{i}\}:=\frac{1}{r}\sum_{j=1}^{r}\sigma_{j}\{z_{i}\}$. We denote by $\mathds{1}\{x\in A\}$ the indicator function, equal to $1$ if $x\in A$ and to 0 if $x\notin A$, for some set $A$. We denote by $\mathrm{SO}(k)$ the group of rotation matrices in $\mathbb{R}^{k\times k}$, given by $\mathrm{SO}(k)\coloneqq\{R\in\mathbb{R}^{k\times k}\mid R^{\top}R=I,\;\det(R)=1\}$. Moreover, we denote by $\mathrm{Diag}(k)$ the space of diagonal matrices in $\mathbb{R}^{k\times k}$. We assume that if the objective function of a minimization (maximization) problem is expressed as the difference of two terms, both evaluating to $\infty$, then the objective function value is interpreted as $\infty$ (-$\infty$). This aligns with the extended arithmetic rules in Rockafellar and Wets [2009].
@@ 2 Minimum-Volume Covering Set
In this section, we consider the general problem of finding the minimum-volume set that contains a fraction $1-\alpha$ of a given set of points $\{y_{1},\dots,y_{n}\}$ in $\mathbb{R}^{k}$, for some $\alpha\in(0,1)$. In Section 2.3, we show how this formulation extends naturally to regression, where we observe pairs $\{(x_{i},y_{i})\}_{i=1}^{n}$ and construct adaptive prediction regions that account for multivariate uncertainty. To define these sets, we consider arbitrary norms $\|\cdot\|$ on $\mathbb{R}^{k}$. Formally, given $\|\cdot\|$, we define:
@@
A notable case arises when $\|\cdot\|$ is the $p$-norm, denoted $\|\cdot\|_{p}$, for some $p\in(0,\infty)$. In this setting, we use the notation $\mathbb{B}(p,M,\mu)$. Beyond optimizing $M$ and $\mu$, we also learn $p$, allowing the norm structure to be selected adaptively based on the data. More generally, combining multiple $p$-norms allows for asymmetric prediction sets, overcoming the symmetry constraints of individual norms. This enables better adaptation to the geometry of the data while ensuring the required $1-\alpha$ coverage. A formal definition of this multi-norm formulation is presented in Section 2.2.
@@ 2.1 Arbitrary norm-based sets
Given a norm $\|\cdot\|$ on $\mathbb{R}^{k}$, our objective is to find the minimum-volume norm-based set that satisfies a prescribed coverage constraint. This leads to the optimization problem:
@@
Here, the constraint ensures that the set contains at least nr+1n-r+1 of the given points $\{y_{1},\dots,y_{n}\}$. The quantity $r$ is chosen such that this corresponds to a fraction $1-\alpha$ of the total points. The volume of a norm-based set is given by
@@
where $\mathbb{B}_{\|\cdot\|}(1)=\{y\in\mathbb{R}^{k}\mid\|y\|\leq 1\}$ is the unit ball under $\|\cdot\|$, and $\lambda$ denotes the Lebesgue measure in $\mathbb{R}^{k}$. Since we do not optimize over the norm $\|\cdot\|$ in this section, the term $\lambda(B_{\|\cdot\|}(1))$ remains fixed, and we can simplify the objective function by minimizing $\log\det(M)$. This reformulation preserves the essential structure of the problem while removing unnecessary constants. We will return to the full volume formulation in Section 2.2, where we optimize over the norm, making the dependence on $\lambda(B_{\|\cdot\|}(1))$ relevant. Under these considerations, the problem reduces to
@@
Throughout this section, we refer to problem (5) as a minimum-volume covering set (MVCS) problem. While the objective function $\log\det(M)$ is convex for $M\succcurlyeq 0$, the problem remains nonconvex due to the cardinality constraint, which introduces a combinatorial structure to the feasible set. This problem is known to be NP hard (see [Ahmadi et al., 2014, Proposition 1]). The following proposition, based on a homogeneity argument, provides an exact reformulation that enables both convex relaxations and exact reformulations via difference-of-convex (DC) programming. In Section 2.2 we also formulate this loss in a way that allows the use of first-order optimization strategies.
@@
Now, due the cardinality constraint, for given $\Lambda$ and $\eta$, the optimal $\nu^{\star}(\Lambda,\eta)$ is equal to the (nr+1)(n-r+1)-th smallest value in the set $\left\{\|\Lambda y_{i}+\eta\|\right\}_{i=1}^{n}$, or, equivalently, the $r$-th largest value, i.e.,
@@
[…]
whose objective function can be rewritten as $k\log\zeta-\log\det(\Lambda)+\zeta\sigma_{r}\left\{\|\Lambda y_{i}+\eta\|\right\}$. Now, since the objective function is convex and differentiable in $\zeta$, we can minimize explicitly over $\zeta$ and obtain
@@
where $\overline{\sigma}_{r}$ and $\overline{\sigma}_{r-1}$ denote the averages of the $r$ and r1r-1 largest values in the set. A formal proof that $f$ and $g$ are convex functions is provided in Appendix A.
A key advantage of the DC reformulation in (9) is that it allows for efficient local optimization using the DC algorithm (DCA) Tao and An [1997]; Horst and Thoai [1999]. This method iteratively constructs convex approximations of the objective function, solving a convex subproblem at each step rather than optimizing the original nonconvex function directly. This approach ensures scalability while preserving the problem’s structure. We briefly outline the DCA approach.
@@
with $\|\cdot\|_{*}$ being the dual norm of $\|\cdot\|$. This definition ensures that $g_{i}$ acts as a supporting hyperplane at the point $\Lambda_{t}y_{i}+\eta_{t}$ with respect to the given norm structure. If the norm is differentiable (which holds for $p>1$ and $x\neq 0$), the subgradient simplifies to:
@@
In the special case of $p=2$, this reduces to $g_{i}={\Lambda_{t}y_{i}+\eta_{t}}/{\|\Lambda_{t}y_{i}+\eta_{t}\|_{2}}$. In practice, problem (10) can be solved efficiently using cvxpy Diamond and Boyd [2016] with the Mosek solver ApS [2024].
A key property of the DCA is that the sequence of objective values $f(\Lambda_{t},\eta_{t})-g(\Lambda_{t},\eta_{t})$ is nonincreasing and convergent (see, e.g., [Niu, 2022, Lemma 1]). However, since problem (6) is nonconvex, DCA does not guarantee global optimality. The algorithm may converge to a local minimum, meaning that the resulting prediction set may not have the smallest possible volume.
@@
Let $(\Lambda^{\star},\eta^{\star})$ denote the final iterate of the DCA method (10) or the global solution of the convex relaxation (11). The corresponding MVCS set, which contains nr+1n-r+1 points, is given by $\{y\in\mathbb{R}^{k}\mid\|\Lambda^{\star}y+\eta^{\star}\|\leq\sigma_{r}\left\{\|\Lambda^{\star}y_{i}+\eta^{\star}\|\right\}\}$. We illustrate in Figure 2 the solutions obtained by both methods for different distributions. In the Gaussian case (left), the solutions are nearly identical, with both approaches aligning with the empirical covariance structure. However, for an asymmetric distribution (right), the convex relaxation method is more influenced by the tails, stretching the prediction set outward, whereas the DCA solution remains concentrated around the high-density region near the origin. These observations suggest that in elliptical distributions, the convex relaxation method performs well, but for non-elliptical distributions, the DCA method provides tighter, more representative sets.
@@
A key insight emerges when taking the derivative of the objective function in (11) with respect to $\nu$: at optimality, $\nu^{\star}$ is the (nr+1)(n-r+1)-th order statistic of the nonconformity scores, enforcing the required coverage constraint. This mechanism closely parallels the structure of quantile regression Koenker [2005], where the pinball loss implicitly selects a quantile level by balancing over- and under-estimation errors. A similar argument appears in [Roth, 2022, Lemma 2.2.1], which establishes that minimizing the pinball loss aligns with quantile estimation through a first-order optimality condition.